OpenStudy (anonymous):

Help! Write the slope-intercept equation of the lines passing through the point (0,-5) and perpendicular to y=x-4

OpenStudy (anonymous):

Okay, so perpendicular means the slope is going to be the opposite reciprocal of the given slope (i.e. if the given slope is 1/4, then the line perpendicular to it will have a slope of -4)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

So if the given line (y=x-4) has a slope of 1 (the coefficient of x), then the slope of the new line will be -1

OpenStudy (anonymous):

So the slope is the coefficient of x?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

If you are given the equation of a line in the form y = ax + b, a is the slope of the line, and b is the intercept (since when you plug in 0 for x, you are left with b)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So whenever they give you a line and tell you to find a line that is perpendicular to it, the slope is going to be the opposite reciprocal, and if they tell you to find a line that is parallel to it, the slopes are the same (going in the same direction)

OpenStudy (anonymous):

so would this form be correct? y-0= 1(x-(-5)

OpenStudy (anonymous):

That's it!

OpenStudy (anonymous):

Then you just multiply it out and get it in the form y = ax + b and you are done

OpenStudy (anonymous):

I got..y-0 =-1(x-(-5) y-0=-x+5 y-o+0=-x+5+0 y=-x+5

OpenStudy (anonymous):

You're almost there, just made a mistake when factoring: y - 0 = -1(x - (-5)) y = -1(x + 5) y = -x - 5

OpenStudy (anonymous):

Wow..thanks a lot

OpenStudy (anonymous):

No problem!